estimate the number of eletrons in 10g of water and  find chargeon them
 
 
 

Grade:6

1 Answers

Arun
6 years ago

no. of moles of water =given mass / molar mass = 10/18 = 5/9 moles

total no. of atoms in 100g of water = 6.022 *1023 * 5/9 = 0.3455 *1024 atoms

no. of electrons in 1 atom of H2O = 2*1 +8 = 10

total no. of elctrons = 0.3455*1024 *10 = 0.3455*1025 electrons..

charge in 1 electron = 1.6 *10-19

charge carried by 0.3455*1025 electrons = 1.6*10-19 * 0.3455*1025 = 5.528 * 105